Estimate each of the following using general rule: 
730 + 998 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

730 + 998
Round off to hundreds
730 is rounds off to 700
998 is rounds off to 1000
Hence, 730 + 998 = 700 + 1000 = 1700
​​
Thanks